OpenStudy (anonymous):

Write the equation of the line perpendicular to the line x - 3y = 9 and passing through the point (3,5). Write the equation of the line parallel to the line 4x + 2y = 5 and passing through the point (-3,5). Write the equation of the line parallel to the line 10x - 5y = 8 and passing through the point (2,4). Write the equation of the line parallel to the line 4x + y = -1 and passing through the point (5,0). Write the equation of the line perpendicular to the line x - 5y = -10 and passing through the point (2,5).

OpenStudy (anonymous):

A) y = -2x - 1 B) y = -4x + 20 C) y = 2x D) y = -3x + 14 E) y = -5x + 15

OpenStudy (anonymous):

for 1st,its d

OpenStudy (anonymous):

what bout the others? :)

OpenStudy (anonymous):

which one to attempt first this or other one??

OpenStudy (anonymous):

I need help with this

OpenStudy (anonymous):

Write the equation of the line parallel to the line 4x + 2y = 5 and passing through the point (-3,5). Write the equation of the line parallel to the line 10x - 5y = 8 and passing through the point (2,4). Write the equation of the line parallel to the line 4x + y = -1 and passing through the point (5,0). Write the equation of the line perpendicular to the line x - 5y = -10 and passing through the point (2,5).

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you need the answers only or explaination......

OpenStudy (anonymous):

answers

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Is the last one E?

OpenStudy (anonymous):

for second one its A

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what about the third

OpenStudy (anonymous):

FOR THIRD ONE ITS C

OpenStudy (anonymous):

FOR FOURTH ONE ITS B

OpenStudy (anonymous):

And the last one E

OpenStudy (anonymous):

YA.......

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

SO .........MY FAN

OpenStudy (anonymous):

heyy :)

OpenStudy (anonymous):

Can you help me with more please im bout to just post it

OpenStudy (anonymous):

I WAS ACTUALLY ASKING FOR THE SAME?????WILL YOU BE?